# A Survey on Training-free Alignment of Large Language Models

### v1

@@ 1 Introduction
VILMO (Duan et al., 2024) , leaf, text width=32.5em ] ] [ Enhanced Prompt
@@
ARL2 Zhang et al. (2024b), , leaf, text width=32.5em ] ] [ Detector-Based [ VLMGUARD  Du et al. (2024), CIDER Xu et al. (2024a),
@@
OPAD Zhu et al. (2025), SCANS Cao et al. (2025), IAR Li et al. (2025) , leaf, text width=32.5em ] ] [ Dynamic Search
@@
However, as these models permeate critical domains such as healthcare, education, and public discourse, their widespread adoption has ignited a dual-edged societal response: urgent concerns over risks and growing demands for enhanced capabilities. On one hand, LLMs bring about various risks like generating harmful content (Soice et al., 2023), perpetuating biases and compromising privacy (Salecha et al., 2024; Staab et al., 2024; Abdulhai et al., 2024). On the other hand, users and industries increasingly demand models that adapt to dynamic knowledge (Zhang et al., 2023; Si et al., 2023; Beukman et al., 2023), excel in specialized tasks (Li et al., 2024a; Chen et al., 2024a), and deliver personalized services (Richardson et al., 2023; Zhuang et al., 2024). These dual imperatives—mitigating risks and fulfilling functional demands—underscore the necessity of aligning LLMs with human values, ethical standards, and practical requirements. Alignment is no longer a technique alone but a prerequisite for responsible and effective deployment.
As a critical problem in artificial intelligence and natural language processing, LLM alignment has attracted considerable research attention. Traditional alignment methods primarily rely on fine-tuning (FT), where model parameters are adjusted using curated datasets. While effective in specific scenarios, these approaches face one or more of the following three critical limitations: (1) Severe Knowledge Degradation: FT may overwrite pretrained knowledge, leading to catastrophic forgetting of general capabilities. (2) Resource Intensity: Collecting high-quality alignment data and retraining LLMs requires prohibitive computational and human resources. (3) Access Constraint: Proprietary models (e.g., GPT-4, Gemini) restrict parameter access, rendering FT infeasible.
@@ 2.1 What are LLM Alignment and TF Alignment?
LLM alignment is the process of ensuring that language models behave in ways that align with human expectations, societal values, legal standards, and ethical principles. This review focuses on TF alignment, which refers to achieving alignment without requiring post-training of the LLM to be aligned111 In this review, base models must be sufficiently capable of responding effectively to TF alignment techniques, which means that these models might require prior fine-tuning..
@@ 3.1 Pre-Decoding TF Alignment
One of the simplest yet effective alignment strategies involves crafting specific prompts to elicit desired responses. ICL has emerged as a powerful tool for aligning LLMs with human preferences, a technique known as In-Context Alignment (ICA). For instance, URIAL (Lin et al., 2023) demonstrates that just three in-context examples and a single system prompt can effectively align pre-trained LLMs, achieving performance comparable to FT methods while significantly reducing costs. Similarly, Anthropological Prompting (AlKhamissi et al., 2024) and VILMO (Duan et al., 2024) leverage specific prompts to enhance alignment. By introducing safety configurations as part of the system prompt, CoSA (Zhang et al., 2024a) enables models to generate outputs that meet specific requirements while accommodating the diverse needs of different cultures, applications, or users.
Although MLLMs introduce a novel visual modality, enabling the embedding of harmful content within images to circumvent security mechanisms, the prompt engineering strategy in pre-decoding remains applicable. AdaShield (Wang et al., 2024c) adds defense prompts before inputs, enhancing safety of MLLMs. It includes both manual static and adaptive defense prompts, the latter being iteratively optimized through collaboration between the target MLLM and an LLM-based defense prompt generator. ICD (Wei et al., 2024), RapGuard (Jiang et al., 2024) and BBA (Zhao et al., 2024) generate scenario-specific security prompts, utilizing multimodal reasoning to identify and mitigate risks dynamically.
@@
Beyond simple prompts, this category encompasses methods that enhance prompts through additional techniques or iterative refinement. OPO (Xu et al., 2023) uses retrieval-augmented-generation (RAG) to address the ever-changing nature of alignment in human values. AUTOCAP (Zhang et al., 2024d) integrates Chain-of-Thought (CoT) reasoning paths across languages to improve cross-lingual alignment. Other methods focus on feeding inputs in a more structured way. For example, Zhang et al. (2024e) introduce goal priority through few-shot prompting, instructing the LLM to prioritize safety over helpfulness. PRETTY (Zhan et al., 2024) improves alignment by adding task-related prior markers to the input prefix, narrowing the performance gap between TF models and FT models. BPO (Cheng et al., 2024) optimizes user prompts to suit LLMs’ input understanding, ensuring user intents are realized without updating internal parameters. PEARL (Fu et al., 2024a) paraphrases questions in expressions preferred by the model, ensuring better alignment with user expectations.
@@
Pre-decoding methods excel in simplicity and are compatible with black box models but face generalization limits due to reliance on few-shot examples and manual prompt design.
@@ 3.2 In-Decoding TF Alignment
For guidance-based approaches, Qian et al. (2024); Wang et al. (2024b); Huang et al. (2024c); Bhattacharjee et al. (2024); Cao et al. (2025); Li et al. (2025) extract guidance vectors to adjust the activation states of the target model during inference, achieving alignment. To predict next-token rewards for efficient and effective autoregressive generation, GenARM Xu et al. (2025) leverages the autoregressive reward model, demonstrating that this parametrization can provably guide frozen LLMs toward any distribution. PAD (Chen et al., 2024b) leverages token-level personalized rewards to guide the decoding process, dynamically tailoring the base model’s predictions to personalized preferences. Nudging Fei et al. (2024) combines a large base model with a much smaller aligned model performs on par with the aligned version of the large base model. RDS Zeng et al. (2025) designs a root classifier based on the discriminative capacity of queries, and then reorders the token and prioritizes the benign token. To directly align model outputs with human preferences, OPAD Zhu et al. (2025) guides the generation of a final output that aligns with the target principle in a token-by-token manner.
@@
Search strategies have also been frequently used for alignment. RAIN (Li et al., 2024b) mirrors human behavioral patterns: contemplating, weighing, and reflecting on the consequences before speaking. It allows LLMs to evaluate their own generation and use the evaluation results to guide rewind and generation for self-alignment. Similarly, DeAL (Huang et al., 2024a) views decoding as a heuristic-guided search process and facilitates the use of a wide variety of alignment objectives. TreeBoN (Qiu et al., 2024) combines tree search strategies with Best-of-N (BoN) sampling, iteratively expanding high-quality partial responses while pruning low-quality candidates to improve alignment quality and efficiency. DARWIN (Hung et al., 2024) achieves the right balance between exploration and exploitation of rewards during decoding with evolutionary heuristics. Additionally, Zhu et al. (2024) proposed PAS, which includes direction search and distance search. They focuse on the alignment with personal traits and develop an activation intervention optimization method to enhance LLMs’ ability to efficiently align with individual behavioral preferences using minimal data and computational resources.
@@
In-decoding TF alignment methods provide dynamic, fine-grained control over model outputs by adjusting hidden states, modifying logits, employing tree search and so on. For MLLMs, pulling hidden states back into the representation space optimized by LLMs are particularly promising, as they address the unique challenges introduced by multimodal integration. While highly effective, this line of research often requires access to model internals, limiting its applicability in black-box systems.
@@ 3.3 Post-Decoding TF Alignment
Post-decoding TF alignment methods are model-agnostic, making them highly versatile. For MLLMs, the integration of visual and textual data provides additional opportunities for dynamic defense mechanisms, though it also increases the complexity of alignment tasks. However, they often introduce latency caused by the filtering or correction steps.
3.4 Numerical Analysis
@@
Results The results are shown in Table 1. Notably, TF alignment methods can match or even exceed the safety and helpfulness performance of FT alignment methods. For example, SCANS outperforms SafeDecoding on SafeEdit and TruthfulQA datasets. It also shows performance comparable to SafeDecoding on AdvBench. This indicates that TF alignment methods warrant further research, particularly in scenarios where FT alignment methods are constrained (e.g., severe knowledge degration). However, TF alignment methods also exhibit certain limitations. For example, TF methods like URIAL and RA-LLM compromise the model’s helpfulness, which certainly damages the helpful performance on TruthfulQA, albeit such damage is less severe compared to the FT method SafeDecoding. More challenges will be described in the next section. We also analyze practical deployment considerations for different TF alignment methods in Table 2 of Appendix A.
@@ 4.1 Open Challenges
While TF alignment offers practical advantages, our systematic analysis reveals three fundamental challenges that constrain its broader adoption: (1) Degradation of General Capabilities: While most TF alignment methods preserve model knowledge integrity better than FT alignment methods, certain in-decoding methods potentially compromise general performance. (2) Increased Inference Overhead: TF alignment approaches often introduce additional latency. Whether through input detection, logits adjustment, or modification of generated responses, the additional computational steps result in higher latency. (3) Difficult to Generalize: TF alignment cannot leverage extensive alignment data. As a result, it may fail to generalize effectively to unseen or more challenging scenarios (e.g. adversarial attacks), as it lacks exposure to diverse or complex cases.
@@ 5 Conclusion
While this survey provides a systematic overview of TF alignment methods for LLMs, several inherent limitations in its scope need more discussion: Firstly, we pay more attention to safety TF alignment due to page limitation, which makes it infeasible to provide a detailed overview of all TF alignment approaches. Thus, we choose to focus intensively on one category. Secondly, experiments just conducted on llama2-7b-chat because the nature of our literature survey. It is hoped that our preliminary exploration will inspire further researches on TF alignment.

### v2

@@ 1 Introduction
ICDPO Song et al. (2025a), VILMO (Duan et al., 2024) , leaf, text width=32.5em ] ] [ Enhanced Prompt
@@
P-Aligner (Song et al., 2025b), ARL2 Zhang et al. (2024b), , leaf, text width=32.5em ] ] [ Detector-Based [ VLMGUARD  Du et al. (2024), CIDER Xu et al. (2024a),
@@
OPAD Zhu et al. (2025), SCANS Cao et al. (2025), IAR Li et al. (2025b),
WSD (Song et al., 2025c), CARDS (Li et al., 2025a),
ARGS (Khanov et al., 2024)
, leaf, text width=32.5em ] ] [ Dynamic Search
@@
However, as these models permeate critical domains such as healthcare, education, and public discourse, their widespread adoption has ignited a dual-edged societal response: urgent concerns over risks and growing demands for enhanced capabilities. On one hand, LLMs bring about various risks like generating harmful content (Soice et al., 2023), perpetuating biases and compromising privacy (Salecha et al., 2024; Staab et al., 2024; Abdulhai et al., 2024). On the other hand, users and industries increasingly demand models that adapt to dynamic knowledge (Zhang et al., 2023; Si et al., 2023; Beukman et al., 2023), excel in specialized tasks (Li et al., 2024a; Chen et al., 2024a), and deliver personalized services (Richardson et al., 2023; Zhuang et al., 2024). These dual imperatives, namely mitigating risks and fulfilling functional demands, underscore the necessity of aligning LLMs with human values, ethical standards, and practical requirements. Alignment is no longer a technique alone but a prerequisite for responsible and effective deployment.
As a critical problem in artificial intelligence and natural language processing, LLM alignment has attracted considerable research attention (Shen et al., 2023; Wang et al., 2024d; Gabriel, 2020; Ouyang et al., 2022). Traditional alignment methods primarily rely on fine-tuning (FT), where model parameters are adjusted using curated datasets. While effective in specific scenarios, these approaches face one or more of the following three critical limitations: (1) Severe Knowledge Degradation: FT may overwrite pretrained knowledge, leading to catastrophic forgetting of general capabilities. (2) Resource Intensity: Collecting high-quality alignment data and retraining LLMs requires prohibitive computational and human resources. (3) Access Constraint: Proprietary models (e.g., GPT-4, Gemini) restrict parameter access, rendering FT infeasible.
@@ 2.1 What are LLM Alignment and TF Alignment?
LLM alignment is the process of ensuring that language models behave in ways that align with human expectations, societal values, legal standards, and ethical principles. This review focuses on TF alignment, which encompasses functional alignment and normative alignment.
Functional alignment emphasizes the model’s ability to perform tasks accurately and reliably. It can be further divided into two categories: (1) Task-specific Alignment: Tailoring the model for specific applications (e.g., Zhan et al. (2024)) to optimize the objective towards particular requirements. (2) General Capability Alignment: Ensuring the model maintains updated knowledge and strong generalization abilities across diverse tasks and scenarios (e.g., Zhang et al. (2024c)).
Normative alignment focuses on ethical, safety, and user-centric aspects. It also contains two main components: (1) Public Safety and Ethical Alignment: Ensuring model outputs comply with safety standards, ethical norms, and regulatory guidelines to avoid harmful content and data privacy violations (e.g., Fang et al. (2024)). The goal is to mitigate potential risks and prevent adverse impacts on society. (2) Personalized Alignment: Adapting the model’s interaction style and outputs to individual preferences and usage patterns, balancing safety protocols with personalized experiences (e.g., Chen et al. (2024b)).
In conclusion, functional alignment ensures the model performs tasks effectively, while normative alignment guarantees that its operation remains safe and ethical.
@@ 3.1 Pre-Decoding TF Alignment
One of the simplest yet effective alignment strategies involves crafting specific prompts to elicit desired responses. ICL has emerged as a powerful tool for aligning LLMs with human preferences, a technique known as In-Context Alignment (ICA). For instance, URIAL (Lin et al., 2023) demonstrates that just three in-context examples and a single system prompt can effectively align pre-trained LLMs, achieving performance comparable to FT methods while significantly reducing costs. Similarly, Anthropological Prompting (AlKhamissi et al., 2024) and VILMO (Duan et al., 2024) leverage specific prompts to enhance alignment. Besides, ICDPO (Song et al., 2025a) leverages ICL and instant scorer to enhance the final performance. To tackle the inflexibility of existing alignment methods when confronted with diverse social norms across different cultures and regions, CoSA (Zhang et al., 2024a) introduces safety configs, which allow models to dynamically adapt during inference based on these configurations, thereby catering to varying requirements.
Although MLLMs introduce a novel visual modality, enabling the embedding of harmful content within images to circumvent security mechanisms, the prompt engineering strategy in pre-decoding remains applicable. To defend against structured-based jailbreak attacks, AdaShield (Wang et al., 2024c) includes both manual static and adaptive defense prompts, the latter being iteratively optimized through collaboration between the target MLLM and an LLM-based defense prompt generator. Similarly, to address the limitation that static safety guidelines fail to account for specific risks in different multimodal contexts, ICD (Wei et al., 2024), RapGuard (Jiang et al., 2024), and BBA (Zhao et al., 2024) generate scenario-specific security prompts. They leverage multimodal reasoning to dynamically identify and mitigate risks.
@@
Beyond simple prompts, this category encompasses methods that enhance prompts through additional techniques or iterative refinement. Certain human values to be aligned often vary with time and place, OPO (Xu et al., 2023) uses retrieval-augmented-generation (RAG) can address the ever-changing nature of alignment in human values. AUTOCAP (Zhang et al., 2024d) integrates Chain-of-Thought (CoT) reasoning paths across languages to improve cross-lingual alignment (Shi et al., 2022; Tanwar et al., 2023; Qin et al., 2023). P-Aligner (Song et al., 2025b) automatically rewrites user instructions with pre-defined principles, significantly boosting downstream LLM alignment. Other methods focus on feeding inputs in a more structured way. For example, Zhang et al. (2024e) introduce goal priority through few-shot prompting, instructing the LLM to prioritize safety over helpfulness. PRETTY (Zhan et al., 2024) improves alignment by adding task-related prior markers to the input prefix, narrowing the performance gap between TF models and FT models. BPO (Cheng et al., 2024) optimizes user prompts to suit LLMs’ input understanding, ensuring user intents are realized without updating internal parameters. PEARL (Fu et al., 2024a) paraphrases questions in expressions preferred by the model, ensuring better alignment with user expectations.
@@
Pre-decoding methods aim to reformulate queries, append exemplars, or leverage detectors before feeding inputs into LLMs, ensuring model outputs align with human standards. These approaches offer advantages of simplicity and broad applicability, working seamlessly with both open-source and closed-source models. However, they suffer from generalization limitations due to reliance on few-shot examples or manual prompt engineering. For instance, CoSA (Zhang et al., 2024a) introduces safety configurations to enable dynamic inferential adaptation, yet this only addresses safety scenarios—cross-scenario generalization, cultural variance, and model-agnostic adaptability remain underexplored. Additionally, most prompt engineering techniques are constrained by context window limits (e.g., LLaMA-2-13B’s 4096-token boundary (Machlab and Battle, 2024)) or suffer semantic drift (Ji et al., 2023) due to excessively long inputs. The subsequent in-decoding and post-decoding methods tackle these limitations from alternative perspectives.
@@ 3.2 In-Decoding TF Alignment
For guidance-based approaches, Qian et al. (2024); Wang et al. (2024b); Huang et al. (2024c); Bhattacharjee et al. (2024); Cao et al. (2025); Li et al. (2025b) extract guidance vectors to adjust the activation states of the target model during inference, achieving alignment. To predict next-token rewards for efficient and effective autoregressive generation, GenARM Xu et al. (2025) and CARDS (Li et al., 2025a) leverage the reward model to guide frozen LLMs toward expected distribution. PAD (Chen et al., 2024b) and ARGS (Khanov et al., 2024) leverage token-level rewards to guide the decoding process, dynamically guiding the base model’s predictions. Nudging Fei et al. (2024) and WSD (Song et al., 2025c) combine a large base model with a much smaller aligned model. RDS Zeng et al. (2025) designs a root classifier based on the discriminative capacity of queries, and then reorders the token and prioritizes the benign token. To directly align model outputs with human preferences, OPAD Zhu et al. (2025) guides the generation of a final output that aligns with the target principle in a token-by-token manner.
@@
Search strategies have also been frequently used for alignment. RAIN (Li et al., 2024b) mirrors human behavioral patterns, which allows LLMs to evaluate their own generation and use the evaluation results to guide rewind and generation for self-alignment. Similarly, DeAL (Huang et al., 2024a) views decoding as a heuristic-guided search process and facilitates the use of a wide variety of alignment objectives. TreeBoN (Qiu et al., 2024) combines tree search strategies with Best-of-N (BoN) sampling, iteratively expanding high-quality partial responses while pruning low-quality candidates to improve alignment quality and efficiency. DARWIN (Hung et al., 2024) achieves the right balance between exploration and exploitation of rewards during decoding with evolutionary heuristics. Additionally, Zhu et al. (2024) proposed PAS, which includes direction search and distance search. They focuse on the alignment with personal traits and develop an activation intervention optimization method to enhance LLMs’ ability to efficiently align with individual behavioral preferences using minimal data and computational resources.
@@
In-decoding TF alignment methods enable dynamic, fine-grained output control via hidden state adjustment, logit modification, tree search strategies, etc. For MLLMs, projecting hidden states into LLM-optimized representation spaces shows promise in resolving multimodal integration challenges. While highly effective for alignment and alleviating certain generalization issues, this paradigm often requires access to model internals, restricting applicability in black-box systems. Moreover, techniques like Trustworthiness and Category-Specific Steering (Qian et al., 2024; Bhattacharjee et al., 2024), which modify hidden or activation states, can inadvertently compromise the model’s pre-existing task-specific knowledge. These trade-offs directly inform the research directions outlined in Section 4.2.
@@ 3.3 Post-Decoding TF Alignment
Analogous to pre-decoding approaches, post-decoding TF alignment methods are model agnostic. Also, they mitigate the generalization constraints seen in pre-decoding stages, offering heightened versatility. In MLLM contexts, integrating visual-textual data enables dynamic defense mechanisms, though this amplifies alignment complexity. A notable trade-off is the latency introduced by filtering or correction processes.
3.4 Quantitative Analysis
@@
Results The results are shown in Table 1. Notably, TF alignment methods can match or even exceed the safety and helpfulness performance of FT alignment methods. For example, SCANS outperforms SafeDecoding on SafeEdit and TruthfulQA datasets. It also shows performance comparable to SafeDecoding on AdvBench. This reflects the effectiveness of TF alignment as a supplement to FT alignment, demonstrating that TF methods merit deeper investigation, particularly in scenarios where FT alignment faces practical constraints (e.g., incompatibility with closed-source models, significant knowledge degradation during fine-tuning). However, TF alignment methods also exhibit certain limitations. For example, TF methods like URIAL and RA-LLM compromise the model’s helpfulness, which certainly damages the helpful performance on TruthfulQA, albeit such damage is less severe compared to the FT method SafeDecoding. More challenges will be described in the next section. We also analyze practical deployment considerations for different TF alignment methods in Table 2 of Appendix A.
@@ 4.1 Open Challenges
While TF alignment offers practical advantages, our systematic analysis reveals three fundamental challenges that constrain its broader adoption: (1) Degradation of General Capabilities: While most TF alignment methods preserve model knowledge integrity better than FT alignment methods, certain in-decoding methods (Qian et al., 2024; Bhattacharjee et al., 2024) potentially compromise general performance. Because these methods adjust the hidden states or activation states of the target model during inference, such operations undermine the model’s inherent knowledge for other tasks. This interference disrupts the pre-trained model’s generalized ability to handle diverse tasks, as the adjusted states may no longer align with the optimal representations learned for cross-task scenarios. Consequently, the model’s performance in non-target tasks might deteriorate, highlighting the trade-off between task-specific adaptation and maintenance of general capabilities in alignment strategies. (2) Increased Inference Overhead: TF alignment approaches often introduce additional latency. Whether through input detection, logits adjustment, or modification of generated responses, the additional computational steps result in higher latency. (3) Difficult to Generalize: TF alignment cannot leverage extensive alignment data. As a result, it may fail to generalize effectively to unseen or more challenging scenarios (e.g. adversarial attacks), as it lacks exposure to diverse or complex cases.
@@ 5 Conclusion
While this survey provides a systematic overview of TF alignment methods for LLMs, several inherent limitations in its scope need more discussion: Firstly, in this review, base models must be sufficiently capable of responding effectively to TF alignment techniques, which means that these models might require prior fine-tuning. Secondly, we pay more attention to safety TF alignment due to page limitation, which makes it infeasible to provide a detailed overview of all TF alignment approaches. Thus, we choose to focus intensively on one category. Finally, experiments just conducted on llama2-7b-chat because the nature of our literature survey. It is hoped that our preliminary exploration will inspire further researches on TF alignment.
@@
Acknowledgments
This work was supported by a grant from the National Natural Science Foundation of China (NSFC) project (No. 62276193), the Key Laboratory of Computing Power Network and Information Security, Ministry of Education under Grant No. 2024ZD027, and the Fundamental Research Funds for the Central Universities, China (No. 2042022dx0001).